Find the value of a
1 = 1/a + 1/6 + 1/8

Answer by sudhanshu_kmr(1152)


1 = 1/a + 1/6 + 1/8
=> 1 - 1/6 -1/8 = 1/a
=> 17/24 = 1/a
=> a = 24/17